OpenStudy (traylea):

Which ordered pair is a solution of the equation y=-11x+4? **WILL MEDAL**

OpenStudy (traylea):

a. (0,-7) b. (-1,-7) c. (1,-7) d. (2,26)

OpenStudy (studygurl14):

Try plugging in each answer choice until you find one that works

OpenStudy (traylea):

@StudyGurl14 that's where I need help I don't understand how

OpenStudy (studygurl14):

Remembered an ordered pair is (x,y) You have y = -11x + 4 Plug in For example, using answer choice A (0,-7) (-7) = -11(0) + 4 -7 \(\neq\) 4 So a is out

OpenStudy (traylea):

I'm sorry I still don't understand how to do it @StudyGurl14

OpenStudy (studygurl14):

Notice answer choice A is (0,-7) right?

OpenStudy (traylea):

Yeah

OpenStudy (studygurl14):

Since ordered pairs are (x,y), this means that answer choice A has x = 0, and y = -7

OpenStudy (studygurl14):

They give you the equation of the line y = -11x + 4

OpenStudy (studygurl14):

For an ordered pair to be a solution, it MUST result in a true equation when plugged in

OpenStudy (studygurl14):

So using answer choice A, x = 0, y = -7 y = -11x + 4 Notice the equation has x and y. Plug in the known vallues (-7) = -11(0) + 4 -7 = 0 + 4 -7\(\neq\)4 Since -7 does not equal 4, you know that (0,-7) is not a point on the line

OpenStudy (traylea):

so...b would be (-7)=-11(-1)+4 -7=-1+4

OpenStudy (traylea):

or am I mixing up the numbers @StudyGurl14

OpenStudy (studygurl14):

no, perfect

OpenStudy (studygurl14):

except -11(-1) isn't -1

OpenStudy (traylea):

ummm would it be 0 like the first equation?

OpenStudy (traylea):

@StudyGurl14

OpenStudy (studygurl14):

No. -11 times -1 = 11

OpenStudy (traylea):

So would it be a positive 1

OpenStudy (traylea):

@StudyGurl14

OpenStudy (traylea):

@StudyGurl14 If you're too busy do you know someone who can help so I can get out of your way (I know it sucks I don't understand I'm sorry.)

OpenStudy (studygurl14):

It would be positive 11

OpenStudy (studygurl14):

I'm so sorry! I help so many people at once that I need tags to remind me. Really, feel free to tag me with every post so I am reminded to come back

OpenStudy (traylea):

so... (-7)=-11(11)+4 -7=-1+4 why in the first equation is it 0 though?

OpenStudy (studygurl14):

No, it is (-7) = -11(-1) + 4 -7 = 11 + 4

OpenStudy (traylea):

Oh

OpenStudy (traylea):

so that's wrong too because 15 does not equal 4

OpenStudy (studygurl14):

exactly! See, you're getting it

OpenStudy (studygurl14):

Well, actually 15 does not equal -7, but yeah

OpenStudy (studygurl14):

-7 \(\neq\) 15

OpenStudy (studygurl14):

Now let's do C (1, -7)

OpenStudy (traylea):

so c would be... (-7) = -11(1) + 4 -7 = -11 + 4

OpenStudy (traylea):

-11+4=-7

OpenStudy (studygurl14):

Yep

OpenStudy (studygurl14):

good job

OpenStudy (studygurl14):

So....does C work?

OpenStudy (traylea):

Yes..? I think

OpenStudy (studygurl14):

yes!

OpenStudy (traylea):

yay thank you so much!

OpenStudy (studygurl14):

anytime :)